Jim can fill a pool carring buckets of water in 30 minutes. Sue can do the same job in 45 minutes, and Tony can do the same job in 1 1/2 hours. How quickly can the three fill the pool together?

I have tried soo many ways, and even my toutor could not remember how to do these problems. Please if any way possible, give a solution and a reasoning of how you got your answer. Thank you.

Answer by stanbon(75887)

Jim can fill a pool carring buckets of water in 30 minutes. Sue can do the same job in 45 minutes, and Tony can do the same job in 1 1/2 hours. How quickly can the three fill the pool together?
---------------------------------
Jim Rate = 1/30 job/min
Sue Rate = 1/45 job/min
Tony Rate= 1/90 job/min
Together Rate = 1/x job min
--------------------------------
EQUATION:
1/30 + 1/45 + 1/90 = 1/x
3x + 2x + x = = 90
6x = 90
x = 15 min (Time for them to do the job together)
======================================
Cheers,
Stan H.